# The theory of implicit operations

### v1

@@ 14. Classes without a Beth companion
The proof of 14.1 can easily be adapted to show that both the variety of semigroups and the variety of commutative semigroups lack a Beth companion. To see this, recall that Isbell’s formulas (see 3.14) also form a dominion base for these varieties (see [71, Thm. 2.3] for semigroups and [69, Thm. 1.1] for commutative semigroups). The changes required for adapting the proof of 14.1 are limited to the following. First, the role of the implicit operation $f$ defined by $\varphi=(x\cdot y\thickapprox 1)\sqcap(y\cdot x\thickapprox 1)$ should be taken over by the implicit operation $g$ defined by
@@
which also defines inverses when they exist. In particular, $g$ coincides with $f$ on $\mathbb{Q}$. Moreover, the monoids $\mathbb{Q}$ and $\bm{C}$ appearing in the proof of 14.1 should be replaced by their semigroup reducts. Lastly, the proof of Claim 14.5 uses the fact that $\bm{A}$ has a neutral element, which need not be the case if $\bm{A}$ is an arbitrary semigroup. This problem can be overcome easily by adding a neutral element to $\bm{A}$ in the proof of that claim.
On the other hand, the implicit operation $g$ becomes extendable when restricted to the quasivariety $\mathsf{CCS}$ of cancellative commutative semigroups. In fact, an argument similar to the one detailed in the proof of 11.9(i) shows that the pp expansion of $\mathsf{CCS}$ induced by $g$ is the Beth companion of $\mathsf{CCS}$ and, moreover, is term equivalent to the variety of Abelian groups (inversion is given by $g$ and the neutral element is rendered as the unary operation $x\cdot g(x)$). ∎
@@
To give a better understanding of the applicability of 14.11, we rely on the next characterization of relatively congruence distributive quasivarieties of Heyting algebras, where $\mathsf{HA}$ stands for the variety of Heyting algebras.
@@
Let $\mathsf{K}$ be a quasivariety of Heyting algebras. Since the variety $\mathsf{HA}$ is congruence distributive by 7.2, from [40, Cor. 1.4] it follows that $\mathsf{K}$ is relatively congruence distributive if and only if $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$. Therefore, it only remains to prove that $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$ if and only if there exists a universal class $\mathsf{M}$ such that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ and $\mathsf{M}\subseteq\mathsf{HA}_{\textsc{fsi}}$.
We first establish the implication from left to right. To this end, assume that $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$. Let $\mathsf{M}=\mathbb{U}(\mathsf{K}_{\textsc{rfsi}})$. The 2.9 yields $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rfsi}})$. It follows that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ because $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{M}\subseteq\mathsf{K}$. It is well known (see, e.g., [85, Prop. A.4.3]) that a Heyting algebra is finitely subdirectly irreducible if and only if its greatest element is join irreducible, a property that can be expressed with a universal sentence. Therefore, $\mathsf{HA}_{\textsc{fsi}}$ is a universal class by 2.1(iii). Then
@@
Thus, $\mathsf{M}$ has the desired properties. For the converse implication, assume that there exists a universal class $\mathsf{M}$ such that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ and $\mathsf{M}\subseteq\mathsf{HA}_{\textsc{fsi}}$. 2.13 implies that $\mathsf{K}_{\textsc{rfsi}}=\mathbb{Q}(\mathsf{M})_{\textsc{rfsi}}\subseteq\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Then $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathbb{U}(\mathsf{M})$ by 2.2. As $\mathsf{M}$ is a universal class contained in $\mathsf{HA}_{\textsc{fsi}}$, we obtain
@@
Before presenting its proof, we first establish a series of consequences of 14.11. As every variety of Heyting algebras is congruence distributive (see 7.2), the following is an immediate corollary of 14.11.
@@
We will use the following consequence of 14.11 to show that infinitely many varieties of Heyting algebras lack a Beth companion.
@@
By 14.11 it is sufficient to show that $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{Q}(\mathsf{K})_{\textsc{rsi}})$. Since $\mathsf{K}$ is a finite set of finite Heyting algebras, $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})\subseteq\mathbb{I}(\mathsf{K})$ (see 2.14). Consequently, from 2.13 it follows that $\mathbb{Q}(\mathsf{K})_{\textsc{rsi}}\subseteq\mathbb{I}\mathbb{S}(\mathsf{K})$. Therefore, if $\bm{A}+\bm{C}_{5}$ embeds properly into a member of $\mathbb{Q}(\mathsf{K})_{\textsc{rsi}}$, then it also embeds properly into a member of $\mathsf{K}$, but this contradicts our hypothesis. Thus, $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{Q}(\mathsf{K})_{\textsc{rsi}})$. ∎
We also obtain an analogue of 14.14 for finitely generated varieties of Heyting algebras.
@@
Suppose that (i) holds. By 14.13 it suffices to show that $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{V}(\mathsf{K})_{\textsc{si}})$. Suppose, with a view to contradiction, that there exists $\bm{D}\in\mathbb{V}(\mathsf{K})_{\textsc{si}}$ into which $\bm{A}+\bm{C}_{5}$ properly embeds. Since $\mathsf{K}$ is a finite set of finite Heyting algebras, $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})\subseteq\mathbb{I}(\mathsf{K})$ (see 2.14). Since $\mathbb{V}(\mathsf{K})$ is congruence distributive by 7.2, from 2.12 it follows that $\mathbb{V}(\mathsf{K})_{\textsc{si}}\subseteq\mathbb{H}\mathbb{S}(\mathsf{K})$. Then $\bm{D}\in\mathbb{H}\mathbb{S}(\mathsf{K})$, and so there exists $\bm{B}\in\mathsf{K}$ such that $\bm{D}\in\mathbb{H}\mathbb{S}(\bm{B})$. Since $\bm{A}+\bm{C}_{5}$ embeds into $\bm{D}$, we have $\bm{A}+\bm{C}_{5}\in\mathbb{I}\mathbb{S}\mathbb{H}\mathbb{S}(\bm{B})$. For every class $\mathsf{M}$ we have $\mathbb{S}\mathbb{H}(\mathsf{M})\subseteq\mathbb{H}\mathbb{S}(\mathsf{M})$ (see, e.g., [21, Lem. II.9.2]) and $\mathbb{I}\mathbb{H}(\mathsf{M})=\mathbb{H}(\mathsf{M})$. Consequently, $\mathbb{I}\mathbb{S}\mathbb{H}\mathbb{S}(\bm{B})=\mathbb{H}\mathbb{S}(\bm{B})$, and hence $\bm{A}+\bm{C}_{5}\in\mathbb{H}\mathbb{S}(\bm{B})$. Then (i) implies $\bm{A}+\bm{C}_{5}\cong\bm{B}$. Recall that $\bm{A}+\bm{C}_{5},\bm{B}\in\mathsf{K}$ and all the members of $\mathsf{K}$ are finite. Therefore, since $\bm{A}+\bm{C}_{5}$ properly embeds into $\bm{D}$ and $\bm{D}\in\mathbb{H}\mathbb{S}(\bm{B})$, we have $\lvert\bm{A}+\bm{C}_{5}\rvert<\lvert\bm{D}\rvert\leqslant\lvert\bm{B}\rvert$. Thus, we reached a contradiction because $\bm{A}+\bm{C}_{5}\cong\bm{B}$.
@@
From [91, Thm. 8.1] it follows that the varieties of Gödel algebras with the strong epimorphism surjectivity property are exactly $\mathsf{GA}$ and $\mathbb{V}(\bm{C}_{n})$ for $n\leqslant 4$. Hence, these varieties are their own Beth companions by 11.9(vi). Let $n\geqslant 5$. As $\bm{C}_{n}\cong\bm{C}_{n-4}+\bm{C}_{5}$, from 14.15 it follows that $\mathbb{V}(\bm{C}_{n})$ lacks a Beth companion. ∎
In order to prove 14.11, we first establish a series of useful results.
@@
Let $\mathsf{K}$ be a quasivariety and $f\in\mathsf{ext}_{pp}(\mathsf{K})$. Then $f^{\bm{A}}$ is total for every $\bm{A}\in\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}$.
Proof. First, assume that $\bm{A}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$. Since $f\in\mathsf{ext}_{pp}(\mathsf{K})$ and $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$, by 8.4 there exists $\bm{B}\in\mathsf{K}_{\textsc{rsi}}$ such that $\bm{A}\leqslant\bm{B}$ and $f^{\bm{B}}$ is total. The maximality of $\bm{A}$ implies that $\bm{A}=\bm{B}$. This shows that $f^{\bm{A}}$ is total for every $\bm{A}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$. Then let $\mathsf{M}=\{\bm{A}\in\mathsf{K}:f^{\bm{A}}\text{ is total}\}$. We have $\mathcal{M}(\mathsf{K}_{\textsc{rsi}})\subseteq\mathsf{M}$. Our goal is to show that $\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}\subseteq\mathsf{M}$. Since $\mathsf{M}\subseteq\mathsf{K}$ and $\mathsf{K}$ is a quasivariety, we have $\mathbb{P}(\mathsf{M})\cap\mathsf{K}=\mathbb{P}(\mathsf{M})$. Therefore, applying 9.7 twice with $\mathbb{O}=\mathbb{P}$ and $\mathbb{O}=\mathbb{H}$, we obtain
@@
We may assume that $\mathsf{K}$ has a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. 14.18 implies that $f^{\bm{A}}$ and $f^{\bm{B}}$ are total for every $f\in\mathcal{F}$. Therefore, $\bm{A}[{L}_{\mathcal{F}}]$ and $\bm{B}[{L}_{\mathcal{F}}]$ are defined and are members of $\mathsf{K}[{L}_{\mathcal{F}}]$. It then follows from 13.28(ii) that
@@
From the definition of a dominion it follows immediately that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})$. To prove the other inclusion, let $b\in B$ and assume that $b\notin\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Then there exist $\bm{C}\in\mathsf{K}$ and homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. The 2.9 implies that there exists $\{\bm{C}_{i}:i\in I\}\subseteq\mathsf{K}_{\textsc{rsi}}$ such that $\bm{C}\leqslant\prod_{i\in I}\bm{C}_{i}$ is a subdirect product. For each $i\in I$ let $p_{i}\colon\bm{C}\to\bm{C}_{i}$ be the restriction of the canonical projection. Since $g(b)\neq h(b)$, there exists $i\in I$ such that $(p_{i}\circ g)(b)\neq(p_{i}\circ h)(b)$. As $(p_{i}\circ g){\upharpoonright}_{A}=p_{i}\circ(g{\upharpoonright}_{A})=p_{i}\circ(h{\upharpoonright}_{A})=(p_{i}\circ h){\upharpoonright}_{A}$, the homomorphisms $p_{i}\circ g,p_{i}\circ h\colon\bm{B}\to\bm{C}_{i}$ witness that $b\notin\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})$. Thus, $\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. ∎
@@
Lastly, from (69) it follows that $\mathsf{ker}(\pi_{2})\subseteq\mathsf{ker}(h)$. As a straightforward consequence of 2.6 (see, e.g., [11, Ex. 1.26.8]), we obtain a homomorphism $g\colon\bm{A}_{2}\to\bm{B}$ such that $h=g\circ\pi_{2}$. ∎
We are now ready to prove 14.11.
@@
Therefore, it only remains to show that $\mathsf{d}_{\mathsf{K}}(\bm{D},\bm{B}\times\bm{B})\neq D$. By 14.20 we have
@@
Since $\langle c_{5},c_{4}\rangle\in(B\times B)-D$, it suffices to show $\langle c_{5},c_{4}\rangle\in\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{D},\bm{B}\times\bm{B})$. Let $g,h\colon\bm{B}\times\bm{B}\to\bm{E}$ be a pair of homomorphisms such that $\bm{E}\in\mathsf{K}_{\textsc{rsi}}$ and $g{\upharpoonright}_{D}=h{\upharpoonright}_{D}$. We need to prove that $g(\langle c_{5},c_{4}\rangle)=h(\langle c_{5},c_{4}\rangle)$. Since $\bm{E}\in\mathsf{K}_{\textsc{rsi}}\subseteq\mathsf{K}_{\textsc{rfsi}}$, 14.21 yields that both $g$ and $h$ factor through a projection. We have two cases: either $g$ and $h$ factor through the same projection or not.
@@
Let $\bm{D}$ be as in 14.22. As $\bm{B}\in\mathsf{K}$, $\bm{D}\leqslant\bm{B}\times\bm{B}$, and $\mathsf{K}$ is a quasivariety, we obtain $\bm{D},\bm{B}\times\bm{B}\in\mathsf{K}$. Since $\bm{B}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$ and $\bm{D}\in\mathbb{H}(\bm{B})$, it follows that $\bm{D},\bm{B}\times\bm{B}\in\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}$. If $\mathsf{K}$ had a Beth companion, then 14.19 would imply that $\mathsf{d}_{\mathsf{K}}(\bm{D},\bm{B}\times\bm{B})=D$, contradicting 14.22. Thus, $\mathsf{K}$ lacks a Beth companion. ∎

### v2

@@ 14. Classes without a Beth companion
The proof of Theorem 14.1 can easily be adapted to show that both the variety of semigroups and the variety of commutative semigroups lack a Beth companion. To see this, recall that Isbell’s formulas (see Example 3.14) also form a dominion base for these varieties (see [71, Thm. 2.3] for semigroups and [69, Thm. 1.1] for commutative semigroups). The changes required for adapting the proof of Theorem 14.1 are limited to the following. First, the role of the implicit operation $f$ defined by $\varphi=(x\cdot y\thickapprox 1)\sqcap(y\cdot x\thickapprox 1)$ should be taken over by the implicit operation $g$ defined by
@@
which also defines inverses when they exist. In particular, $g$ coincides with $f$ on $\mathbb{Q}$. Moreover, the monoids $\mathbb{Q}$ and $\bm{C}$ appearing in the proof of Theorem 14.1 should be replaced by their semigroup reducts. Lastly, the proof of Claim 14.5 uses the fact that $\bm{A}$ has a neutral element, which need not be the case if $\bm{A}$ is an arbitrary semigroup. This problem can be overcome easily by adding a neutral element to $\bm{A}$ in the proof of that claim.
On the other hand, the implicit operation $g$ becomes extendable when restricted to the quasivariety $\mathsf{CCS}$ of cancellative commutative semigroups. In fact, an argument similar to the one detailed in the proof of Theorem 11.9(i) shows that the pp expansion of $\mathsf{CCS}$ induced by $g$ is the Beth companion of $\mathsf{CCS}$ and, moreover, is term equivalent to the variety of Abelian groups (inversion is given by $g$ and the neutral element is rendered as the unary operation $x\cdot g(x)$). ∎
@@
To give a better understanding of the applicability of Theorem 14.11, we rely on the next characterization of relatively congruence distributive quasivarieties of Heyting algebras, where $\mathsf{HA}$ stands for the variety of Heyting algebras.
@@
Let $\mathsf{K}$ be a quasivariety of Heyting algebras. Since the variety $\mathsf{HA}$ is congruence distributive by Theorem 7.2, from [40, Cor. 1.4] it follows that $\mathsf{K}$ is relatively congruence distributive if and only if $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$. Therefore, it only remains to prove that $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$ if and only if there exists a universal class $\mathsf{M}$ such that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ and $\mathsf{M}\subseteq\mathsf{HA}_{\textsc{fsi}}$.
We first establish the implication from left to right. To this end, assume that $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{HA}_{\textsc{fsi}}$. Let $\mathsf{M}=\mathbb{U}(\mathsf{K}_{\textsc{rfsi}})$. The Subdirect Decomposition Theorem 2.9 yields $\mathsf{K}=\mathbb{Q}(\mathsf{K}_{\textsc{rfsi}})$. It follows that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ because $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathsf{M}\subseteq\mathsf{K}$. It is well known (see, e.g., [85, Prop. A.4.3]) that a Heyting algebra is finitely subdirectly irreducible if and only if its greatest element is join irreducible, a property that can be expressed with a universal sentence. Therefore, $\mathsf{HA}_{\textsc{fsi}}$ is a universal class by Theorem 2.1(iii). Then
@@
Thus, $\mathsf{M}$ has the desired properties. For the converse implication, assume that there exists a universal class $\mathsf{M}$ such that $\mathsf{K}=\mathbb{Q}(\mathsf{M})$ and $\mathsf{M}\subseteq\mathsf{HA}_{\textsc{fsi}}$. Theorem 2.13 implies that $\mathsf{K}_{\textsc{rfsi}}=\mathbb{Q}(\mathsf{M})_{\textsc{rfsi}}\subseteq\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{M})$. Then $\mathsf{K}_{\textsc{rfsi}}\subseteq\mathbb{U}(\mathsf{M})$ by Theorem 2.2. As $\mathsf{M}$ is a universal class contained in $\mathsf{HA}_{\textsc{fsi}}$, we obtain
@@
Before presenting its proof, we first establish a series of consequences of Theorem 14.11. As every variety of Heyting algebras is congruence distributive (see Theorem 7.2), the following is an immediate corollary of Theorem 14.11.
@@
We will use the following consequence of Theorem 14.11 to show that infinitely many varieties of Heyting algebras lack a Beth companion.
@@
By Theorem 14.11 it is sufficient to show that $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{Q}(\mathsf{K})_{\textsc{rsi}})$. Since $\mathsf{K}$ is a finite set of finite Heyting algebras, $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})\subseteq\mathbb{I}(\mathsf{K})$ (see Proposition 2.14). Consequently, from Theorem 2.13 it follows that $\mathbb{Q}(\mathsf{K})_{\textsc{rsi}}\subseteq\mathbb{I}\mathbb{S}(\mathsf{K})$. Therefore, if $\bm{A}+\bm{C}_{5}$ embeds properly into a member of $\mathbb{Q}(\mathsf{K})_{\textsc{rsi}}$, then it also embeds properly into a member of $\mathsf{K}$, but this contradicts our hypothesis. Thus, $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{Q}(\mathsf{K})_{\textsc{rsi}})$. ∎
We also obtain an analogue of Corollary 14.14 for finitely generated varieties of Heyting algebras.
@@
Suppose that (i) holds. By Corollary 14.13 it suffices to show that $\bm{A}+\bm{C}_{5}\in\mathcal{M}(\mathbb{V}(\mathsf{K})_{\textsc{si}})$. Suppose, with a view to contradiction, that there exists $\bm{D}\in\mathbb{V}(\mathsf{K})_{\textsc{si}}$ into which $\bm{A}+\bm{C}_{5}$ properly embeds. Since $\mathsf{K}$ is a finite set of finite Heyting algebras, $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})\subseteq\mathbb{I}(\mathsf{K})$ (see Proposition 2.14). Since $\mathbb{V}(\mathsf{K})$ is congruence distributive by Theorem 7.2, from 2.12 it follows that $\mathbb{V}(\mathsf{K})_{\textsc{si}}\subseteq\mathbb{H}\mathbb{S}(\mathsf{K})$. Then $\bm{D}\in\mathbb{H}\mathbb{S}(\mathsf{K})$, and so there exists $\bm{B}\in\mathsf{K}$ such that $\bm{D}\in\mathbb{H}\mathbb{S}(\bm{B})$. Since $\bm{A}+\bm{C}_{5}$ embeds into $\bm{D}$, we have $\bm{A}+\bm{C}_{5}\in\mathbb{I}\mathbb{S}\mathbb{H}\mathbb{S}(\bm{B})$. For every class $\mathsf{M}$ we have $\mathbb{S}\mathbb{H}(\mathsf{M})\subseteq\mathbb{H}\mathbb{S}(\mathsf{M})$ (see, e.g., [21, Lem. II.9.2]) and $\mathbb{I}\mathbb{H}(\mathsf{M})=\mathbb{H}(\mathsf{M})$. Consequently, $\mathbb{I}\mathbb{S}\mathbb{H}\mathbb{S}(\bm{B})=\mathbb{H}\mathbb{S}(\bm{B})$, and hence $\bm{A}+\bm{C}_{5}\in\mathbb{H}\mathbb{S}(\bm{B})$. Then (i) implies $\bm{A}+\bm{C}_{5}\cong\bm{B}$. Recall that $\bm{A}+\bm{C}_{5},\bm{B}\in\mathsf{K}$ and all the members of $\mathsf{K}$ are finite. Therefore, since $\bm{A}+\bm{C}_{5}$ properly embeds into $\bm{D}$ and $\bm{D}\in\mathbb{H}\mathbb{S}(\bm{B})$, we have $\lvert\bm{A}+\bm{C}_{5}\rvert<\lvert\bm{D}\rvert\leqslant\lvert\bm{B}\rvert$. Thus, we reached a contradiction because $\bm{A}+\bm{C}_{5}\cong\bm{B}$.
@@
From [91, Thm. 8.1] it follows that the varieties of Gödel algebras with the strong epimorphism surjectivity property are exactly $\mathsf{GA}$ and $\mathbb{V}(\bm{C}_{n})$ for $n\leqslant 4$. Hence, these varieties are their own Beth companions by Theorem 11.9(vi). Let $n\geqslant 5$. As $\bm{C}_{n}\cong\bm{C}_{n-4}+\bm{C}_{5}$, from Corollary 14.15 it follows that $\mathbb{V}(\bm{C}_{n})$ lacks a Beth companion. ∎
In order to prove Theorem 14.11, we first establish a series of useful results.
@@
Let $\mathsf{K}$ be a quasivariety and $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Then $f^{\bm{A}}$ is total for every $\bm{A}\in\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}$.
Proof. First, assume that $\bm{A}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$. Since $f\in\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and $\bm{A}\in\mathsf{K}_{\textsc{rsi}}$, by Theorem 8.4 there exists $\bm{B}\in\mathsf{K}_{\textsc{rsi}}$ such that $\bm{A}\leqslant\bm{B}$ and $f^{\bm{B}}$ is total. The maximality of $\bm{A}$ implies that $\bm{A}=\bm{B}$. This shows that $f^{\bm{A}}$ is total for every $\bm{A}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$. Then let $\mathsf{M}=\{\bm{A}\in\mathsf{K}:f^{\bm{A}}\text{ is total}\}$. We have $\mathcal{M}(\mathsf{K}_{\textsc{rsi}})\subseteq\mathsf{M}$. Our goal is to show that $\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}\subseteq\mathsf{M}$. Since $\mathsf{M}\subseteq\mathsf{K}$ and $\mathsf{K}$ is a quasivariety, we have $\mathbb{P}(\mathsf{M})\cap\mathsf{K}=\mathbb{P}(\mathsf{M})$. Therefore, applying Proposition 9.7 twice with $\mathbb{O}=\mathbb{P}$ and $\mathbb{O}=\mathbb{H}$, we obtain
@@
We may assume that $\mathsf{K}$ has a Beth companion of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. Proposition 14.18 implies that $f^{\bm{A}}$ and $f^{\bm{B}}$ are total for every $f\in\mathcal{F}$. Therefore, $\bm{A}[{L}_{\mathcal{F}}]$ and $\bm{B}[{L}_{\mathcal{F}}]$ are defined and are members of $\mathsf{K}[{L}_{\mathcal{F}}]$. It then follows from Proposition 13.28(ii) that
@@
From the definition of a dominion it follows immediately that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})$. To prove the other inclusion, let $b\in B$ and assume that $b\notin\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Then there exist $\bm{C}\in\mathsf{K}$ and homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. The Subdirect Decomposition Theorem 2.9 implies that there exists $\{\bm{C}_{i}:i\in I\}\subseteq\mathsf{K}_{\textsc{rsi}}$ such that $\bm{C}\leqslant\prod_{i\in I}\bm{C}_{i}$ is a subdirect product. For each $i\in I$ let $p_{i}\colon\bm{C}\to\bm{C}_{i}$ be the restriction of the canonical projection. Since $g(b)\neq h(b)$, there exists $i\in I$ such that $(p_{i}\circ g)(b)\neq(p_{i}\circ h)(b)$. As $(p_{i}\circ g){\upharpoonright}_{A}=p_{i}\circ(g{\upharpoonright}_{A})=p_{i}\circ(h{\upharpoonright}_{A})=(p_{i}\circ h){\upharpoonright}_{A}$, the homomorphisms $p_{i}\circ g,p_{i}\circ h\colon\bm{B}\to\bm{C}_{i}$ witness that $b\notin\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})$. Thus, $\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. ∎
@@
Lastly, from (69) it follows that $\mathsf{ker}(\pi_{2})\subseteq\mathsf{ker}(h)$. As a straightforward consequence of Proposition 2.6 (see, e.g., [11, Ex. 1.26.8]), we obtain a homomorphism $g\colon\bm{A}_{2}\to\bm{B}$ such that $h=g\circ\pi_{2}$. ∎
We are now ready to prove Theorem 14.11.
@@
Therefore, it only remains to show that $\mathsf{d}_{\mathsf{K}}(\bm{D},\bm{B}\times\bm{B})\neq D$. By Proposition 14.20 we have
@@
Since $\langle c_{5},c_{4}\rangle\in(B\times B)-D$, it suffices to show $\langle c_{5},c_{4}\rangle\in\mathsf{d}_{\mathsf{K}_{\textsc{rsi}}}(\bm{D},\bm{B}\times\bm{B})$. Let $g,h\colon\bm{B}\times\bm{B}\to\bm{E}$ be a pair of homomorphisms such that $\bm{E}\in\mathsf{K}_{\textsc{rsi}}$ and $g{\upharpoonright}_{D}=h{\upharpoonright}_{D}$. We need to prove that $g(\langle c_{5},c_{4}\rangle)=h(\langle c_{5},c_{4}\rangle)$. Since $\bm{E}\in\mathsf{K}_{\textsc{rsi}}\subseteq\mathsf{K}_{\textsc{rfsi}}$, Proposition 14.21 yields that both $g$ and $h$ factor through a projection. We have two cases: either $g$ and $h$ factor through the same projection or not.
@@
Let $\bm{D}$ be as in Claim 14.22. As $\bm{B}\in\mathsf{K}$, $\bm{D}\leqslant\bm{B}\times\bm{B}$, and $\mathsf{K}$ is a quasivariety, we obtain $\bm{D},\bm{B}\times\bm{B}\in\mathsf{K}$. Since $\bm{B}\in\mathcal{M}(\mathsf{K}_{\textsc{rsi}})$ and $\bm{D}\in\mathbb{H}(\bm{B})$, it follows that $\bm{D},\bm{B}\times\bm{B}\in\mathbb{H}\mathbb{P}(\mathcal{M}(\mathsf{K}_{\textsc{rsi}}))\cap\mathsf{K}$. If $\mathsf{K}$ had a Beth companion, then Proposition 14.19 would imply that $\mathsf{d}_{\mathsf{K}}(\bm{D},\bm{B}\times\bm{B})=D$, contradicting Claim 14.22. Thus, $\mathsf{K}$ lacks a Beth companion. ∎